Dear Sirs,
I have a Q, I tried many times but it is still not right. Given that X is a random variable for a normal distribution with mean μ = 54.3 and standard deviation σ = 8.9 find P(X > 35.076).
Please help
Regards
Thong

Answer by rothauserc(4718)

We need to calculate the z-value for P(X < 35.076) and its associated probability by consulting the z-table.
note that P(x > 35.076) = 1 - P(X < 35.076)
z-value = X - mean / std deviation = (35.076 - 54.3) / 8.9 = −2.16
consult table of negative z-values for P(X < 35.076) = 0.0154
P(x > 35.076) = 1 - 0.0154 = 0.9846